# Nudging Nutrition: Lessons from the Danish “Fat Tax”

### v2

@@ 7 Conclusion
The present study has been the first to estimate the causal effect of the Danish tax targeting saturated fats on consumer behavior and expenditure, employing a valid and culturally-close geographic control group consisting of Northern-German households, located in Schleswig-Holstein and Hamburg. Our analysis is based on Callaway and Sant’Anna, ’s (2021) doubly robust DID estimators that allow to flexibly include observed covariates and provide confidence intervals that account for multiple hypothesis testing. The longitudinal nature of our data allowed us to gain valuable insights into the evolution of causal effects over time. To ensure the validity of our findings, we studied the plausibility of the identifying assumptions, most crucially the parallel trends and no anticipation assumption, and performed robustness checks with respect to cross-border shopping.
Distinguishing itself as the first to investigate expenditure patterns, tax abolishment effects, and sustained habit formation post-tax, our research contributes a comprehensive analysis of the policy’s implications. Findings on the purchased weight, the number of packages purchased, and the expenditure paint an interesting picture of the effect of the tax on consumer behavior. Perhaps the most relevant finding is that the tax achieved a reduction in consumption of bacon, cheese, and liver sausage, and salami ranging between 11% and 13%. For bacon, cheese, and liver sausage, the effect of the tax was mainly driven by low-income consumers. This indicates that the tax aligned with its goal of reducing the intake of saturated fats from these products in the population that might benefit most from a reduction. It is important to emphasize, however, that it is unclear whether these products are substituted by healthier alternatives, offering an important avenue for future research. After the tax was repealed, these effects are no longer statistically significant.
The second most significant finding to emerge from this study is the unchanged consumption of margarine, butter, sour cream, and cream, while expenditure on these products increased between 14% and 25%. Interestingly, the increase in expenditure for these products is similar for low and higher-income households, which implies the tax’s regressivity with respect to these products. There are various reasons for this observation, and to improve the design of fat taxes in future, more research is needed to understand why low-income consumers do not substitute away from these products. Exploring the mechanisms behind observed consumer responses is a promising avenue for future research. For instance, gaining information on final consumer characteristics, rather than household characteristics, could provide deeper insights.
Previous studies exclusively focused on the tax’s implementation phase, neglecting post-tax dynamics. Our study, however, sheds light on intriguing post-tax patterns. Notably, the removal of the tax led to varied effects across different product categories. An important finding showed elevated quantities consumed of products high in saturated fat content, even surpassing pre-tax levels in some cases. This raises concerns that the process of implementing and abolishing the tax might have exacerbated unfavorable consumer behavior. For instance, we did observe persisting elevated consumption patterns for butter, margarine, and sour cream in the two years post-tax, indicating that there might be some tax-induced habit formation with respect to this product, especially for high-income households. In future research, it could be interesting to further investigate the longer-term impact (i.e., the impact beyond the year 2014) of the tax’s abolishment.
Our examination of price dynamics has revealed that not all price changes are passed on to the consumers and reflected in the store prices. Among certain product categories (butter, cream, margarine, and sour cream), we noted price increases, while other products displayed no significant price alterations. The abolishment of the tax again resulted in divergent price changes. While prices decreased for some products, such as margarine and liver sausage, others, like butter and sour cream, exhibited persistent elevated prices.
Understanding these effects and the underlying mechanisms is imperative for policymakers. A food tax can be used as a tool to improve public health by influencing consumers’ behavior while increasing revenue from tax. Taken together, our findings do not support strong recommendations to impose a broad tax on saturated fat, as the patterns in consumer responses vary heavily across products. To optimize the design of a tax on saturated fats, more research is required to understand the mechanisms driving consumer choices and whether substitutions towards healthier alternatives are made. Equally important is the consideration of the policy’s impact on low-income households, as our results clearly show that they are disproportionately affected by the tax.
At present, several questions persist regarding the impact of the fat tax. While our analysis considered specific aggregated food categories, an intriguing avenue for future research involves studying substitution dynamics within these product categories. Unfortunately, the study’s scope was constrained by the lack of detailed data on the fat percentages of products within these categories. Access to data on fat percentages could also enable a direct comparison between observed price changes and anticipated adjustments if the tax were entirely shifted to consumers, providing a more thorough analysis of the tax’s pricing dynamics. Furthermore, more research is required to determine the implications of the tax on health outcomes. An extended investigation into health-related consequences, such as potential effects on BMI, would offer a more comprehensive understanding of the tax’s potential long-term health implications. Finally, understanding the impact of health messaging induced by the fat tax and how the public debate shaped attitudes toward products high in saturated fat is a crucial aspect that warrants further study.
@@ A.2 Summary statistics of pre-tax household characteristics
Geographic region Capital region 575 21.3 0 0.0 575 12.8
Zealand 606 22.5 0 0.0 606 13.5
Funen 276 10.2 0 0.0 276 6.2
North Jutland 382 14.2 0 0.0 382 8.5
East Jutland 456 16.9 0 0.0 456 10.2
South/West Jutland 402 14.9 0 0.0 402 9.0
Hamburg 0 0.0 661 37.0 661 14.8
Schleswig-Holstein 0 0.0 1,123 63.0 1,123 25.1
Age (head of hh.) Under 30 573 21.2 202 11.3 775 17.3
30-39 713 26.4 362 20.3 1,075 24.0
40-49 609 22.6 468 26.2 1,077 24.0
50-59 512 19.0 333 18.7 845 18.9
60-69 136 5.0 304 17.0 440 9.8
70 or over 154 5.7 115 6.4 269 6.0
Children under 15 1 2,131 79.0 1,300 72.9 3,431 76.6
2 253 9.4 280 15.7 533 11.9
3 224 8.3 154 8.6 378 8.4
@@
Gender (diary keeper) Male 599 22.2 737 41.3 1,336 29.8
Female 2,098 77.8 1,047 58.7 3,145 70.2
Education 2 543 20.1 413 23.1 956 21.3
(ISCED level) 3 1,011 37.5 666 37.3 1,677 37.4
5 392 14.5 175 9.8 567 12.6
6 578 21.4 211 11.8 789 17.6
7 173 6.4 319 17.9 492 11.0
Income (head of hh.) Very low income 837 31.0 608 34.1 1,445 32.2
Low income 547 20.3 344 19.3 891 19.9
Medium income 361 13.4 298 16.7 659 14.7
High income 451 16.7 214 12.0 665 14.8
Very high income 501 18.6 320 17.9 821 18.3
Household size 1 981 36.4 455 25.5 1,436 32.0
2 1,045 38.8 689 38.6 1,734 38.7
3 274 10.2 327 18.3 601 13.4
4 286 10.6 226 12.7 512 11.4
@@
Country Denmark 2,697 100.0 0 0.0 2,697 60.2
Germany 0 0.0 1,784 100.0 1,784 39.8
all 2,697 100.0 1,784 100.0 4,481 100.0
@@ A.3 Summary statistics for the consumption and price outcomes during the pre-tax and tax period
Product Outcome Pre-tax DK Tax DK $\Delta_{\text{DK}}$ Pre-tax DE Tax DE $\Delta_{\text{DE}}$ Bacon Weight 562.67 486.51 -76.16 382.82 371.16 -11.66 Amount 1.89 1.80 -0.09 1.71 1.68 -0.03 Expenditure 359.84 331.98 -27.86 268.52 274.60 6.08 Price 67.74 71 3.26 79.93 85.08 5.15 Butter Weight 1,607.62 1,554.70 -52.92 1,507.32 1,483.01 -24.31 Amount 6.27 6.22 -0.05 6.13 6.00 -0.13 Expenditure 913.97 1,068.94 154.97 595.71 625.36 29.65 Price 59.65 72.36 12.71 41.06 43.41 2.35 Cheese Weight 2,717.23 2,146.68 -570.55 2,678.16 2,569.77 -108.39 Amount 4.24 3.76 -0.48 9.44 9.30 -0.14 Expenditure 2,306.91 1,912.95 -393.96 1,942.41 1,948.03 5.62 Price 90.50 95.79 5.29 74.37 78.13 3.76 Cream Weight 1,465.81 1,548.64 82.83 1,099.53 1,108.92 9.39 Amount 3.47 3.53 0.06 5.10 5.14 0.04 Expenditure 481.91 569.35 87.44 242.29 267.57 25.28 Price 34.95 39.24 4.29 23.03 25.14 2.11 Liver sausage Weight 1,138.17 977.71 -160.46 538.43 549.05 10.62 Amount 3.31 2.84 -0.47 2.87 2.89 0.02 Expenditure 613.31 530.67 -82.64 374.37 402.22 27.85 Price 61.47 63.19 1.72 75.21 79.22 4.01 Margarine Weight 1,842.00 1,693.05 -148.95 1,969.10 1,760.18 -208.92 Amount 3.87 3.63 -0.24 4.12 3.58 -0.54 Expenditure 511.17 570.55 59.38 368.39 350.34 -18.05 Price 29.35 34.78 5.43 19.87 21.29 1.42 Salami Weight 568.87 480.48 -88.39 490.46 477.05 -13.41 Amount 2.69 2.47 -0.22 3.12 3.02 -0.10 Expenditure 598.54 506.25 -92.29 465.39 466.93 1.54 Price 121.07 121.47 0.40 111.13 113.61 2.48 Sour cream Weight 613.76 604.32 -9.44 506.46 524.83 18.37 Amount 1.33 1.30 -0.03 2.47 2.55 0.08 Expenditure 176.17 186.28 10.11 127.57 132.27 4.70 Price 30.57 32.43 1.86 26.01 25.96 -0.05 DK indicates Danish households and DE indicates Northern-German households. $\Delta$ indicates the difference.
@@ Appendix B Fitted propensity scores for overlap verification
Figure 14: Bacon.
Figure 15: Butter.
Figure 16: Cheese.
Figure 17: Cream.
Figure 18: Liver sausage.
Figure 19: Margarine.
Figure 20: Salami.
@@
Appendix C Evolution of tax effects over time
C.1 Bacon

### v3

@@ 8 Conclusion
This study is the first to causally estimate the impact of Denmark’s saturated fat tax on consumer behavior and expenditure, utilizing a culturally similar control group of Northern German households from Schleswig-Holstein and Hamburg. Employing the doubly robust difference-in-differences estimator by Callaway and Sant’Anna, (2021), we flexibly incorporated observed covariates and accounted for multiple hypothesis testing, enhancing the validity of our results. The longitudinal nature of our data provided valuable insights into the evolution of the tax’s effects over time. We ensured the robustness of our findings by thoroughly examining the plausibility of identifying assumptions, particularly the parallel trends and no anticipation assumptions, and by conducting robustness checks related to cross-border shopping.
Our analysis offers a comprehensive examination of the policy’s implications, including expenditure patterns, the effects of tax abolishment, and post-tax habit formation—areas previously unexplored in the literature. We observed that the fat tax effectively reduced consumption of certain high-fat products, such as bacon, cheese, liver sausage, and salami, with reductions ranging between 11% and 13%. Notably, the reduction in consumption of bacon, cheese, and liver sausage was primarily driven by low-income consumers, indicating that the tax may have effectively targeted the population most in need of dietary improvements. However, it remains unclear whether consumers substituted these products with healthier alternatives, highlighting an important avenue for future research. Importantly, the consumption reductions were not sustained after the tax was repealed, suggesting a lack of long-term behavioral change.
Conversely, we found that consumption of products like margarine, butter, sour cream, and cream remained largely unchanged during the tax period, despite significant increases in expenditure ranging from 14% to 25%. This suggests inelastic demand for these items, with consumers, regardless of income level, absorbing the higher costs without reducing consumption. This uniform increase across income groups indicates a regressive effect of the tax, disproportionately affecting lower-income households. Understanding why low-income consumers did not substitute away from these products is critical for refining the design of such taxes. Future research could delve deeper into consumer behavior mechanisms, potentially through more granular data on individual consumer characteristics.
An important finding of our study is the significant increase in cross-border shopping of butter by Danish households living close to the German border during the tax period. Despite this increase in cross-border purchases, the overall consumption of butter did not decrease significantly, suggesting that cross-border shopping mitigated the intended effect of the tax on reducing saturated fat intake. The ability of consumers to avoid the tax by purchasing goods abroad underscores a critical limitation of the policy, particularly for small countries with accessible borders. Our analysis indicates that the tax’s effectiveness was compromised in regions where cross-border shopping was feasible, highlighting the need for policymakers to consider geographical factors and potential avoidance behaviors when designing such interventions.
Moreover, we observed that the increase in cross-border shopping persisted even after the tax was abolished, indicating a lasting change in consumer behavior. This suggests that once consumers became accustomed to purchasing goods across the border, they continued to do so, further diminishing the potential long-term impact of the tax. Understanding the factors that encourage or discourage cross-border shopping is essential for developing more effective tax policies that achieve public health objectives without unintended consequences.
Our investigation into post-tax dynamics revealed that the abolishment of the fat tax led to varied effects across different product categories. For some products, such as butter, margarine, and sour cream, consumption not only rebounded but exceeded pre-tax levels, particularly among high-income households. This raises concerns that the temporary nature of the tax may have inadvertently encouraged higher consumption of saturated fats after its repeal, possibly due to changes in consumer habits or perceptions. Further research extending beyond 2014 would be valuable to assess the long-term impact of the tax’s removal.
Analysis of price dynamics indicated that the tax was not uniformly passed through to consumers across all products. While prices increased for butter, cream, margarine, and sour cream during the tax period, other products showed no significant price changes. After the tax was abolished, prices decreased for some items like margarine and liver sausage but remained elevated for butter and sour cream. These findings suggest complexities in how taxes are transmitted through the supply chain and underscore the need for policymakers to consider market structures and pricing behaviors when designing such taxes.
Overall, our findings suggest that a broad tax on saturated fats may not uniformly achieve desired public health outcomes, given the varied consumer responses across different products and the mitigating effect of cross-border shopping. The tax’s regressive impact on low-income households further complicates its efficacy as a policy tool. To optimize the design of fat taxes, it is essential to understand the underlying mechanisms driving consumer choices, consider geographical and cross-border factors, and to implement targeted approaches that minimize adverse effects on vulnerable populations. Policymakers should also contemplate complementary measures, such as international cooperation on tax policies, subsidies for healthier alternatives, or educational campaigns, to encourage lasting dietary changes.
Several limitations of our study point to areas for future research. Our analysis was constrained by the lack of detailed data on the fat content of individual products, limiting our ability to assess substitution dynamics within product categories or to compare observed price changes with theoretical expectations of full tax pass-through. Additionally, we did not examine the direct health outcomes associated with the tax, such as changes in body mass index (BMI) or cardiovascular health indicators. Investigating these health-related consequences would provide a more comprehensive understanding of the tax’s long-term implications. Lastly, exploring the role of health messaging and public perception during the tax’s implementation and repeal could shed light on the sociocultural factors influencing consumer behavior.
In conclusion, while the Danish fat tax had some success in reducing consumption of specific high-fat products among certain consumer groups, its effectiveness was undermined by cross-border shopping and varied responses across products and income levels. The ability of consumers to avoid the tax by purchasing goods abroad highlights a significant challenge in implementing such policies, especially in countries with accessible borders. Future policies should consider these nuances to design more effective strategies that promote healthier eating habits without disproportionately impacting low-income households or being circumvented through cross-border purchases.
@@ A.2 Summary statistics of pre-tax household characteristics
Geographic region Capital region 583 21.5 0 0.0 583 12.9
Zealand 606 22.3 0 0.0 606 13.5
Funen 276 10.2 0 0.0 276 6.1
North Jutland 385 14.2 0 0.0 385 8.6
East Jutland 461 17.0 0 0.0 461 10.2
South/West Jutland 404 14.9 0 0.0 404 9.0
Hamburg 0 0.0 661 37.0 661 14.7
Schleswig-Holstein 0 0.0 1,125 63.0 1,125 25.0
Age (head of hh.) Under 30 578 21.3 203 11.4 781 17.4
30-39 717 26.4 362 20.3 1,079 24.0
40-49 614 22.6 469 26.3 1,083 24.1
50-59 514 18.9 333 18.6 847 18.8
60-69 137 5.0 304 17.0 441 9.8
70 or over 155 5.7 115 6.4 270 6.0
Children under 15 2,145 79.0 1,302 72.9 3,447 76.6
2 254 9.4 280 15.7 534 11.9
3 227 8.4 154 8.6 381 8.5
@@
Gender (diary keeper) Male 606 22.3 738 41.3 1344 29.9
Female 2,109 77.7 1,048 58.7 3,157 70.1
Education 2 547 20.1 413 23.1 960 21.3
(ISCED level) 3 1,015 37.4 666 37.3 1,681 37.4
5 396 14.6 176 9.8 572 12.7
6 583 21.5 211 11.8 794 17.6
7 174 6.4 320 17.9 494 11.0
Income (head of hh.) Very low income 843 31.1 610 34.1 1453 32.3
Low income 552 20.3 344 19.3 896 19.9
Medium income 362 13.3 298 16.7 660 14.7
High income 455 16.8 214 12.0 669 14.9
Very high income 503 18.5 320 17.9 823 18.3
Household size 1 990 36.5 457 25.6 1447 32.1
2 1,051 38.7 689 38.6 1,740 38.7
3 275 10.1 327 18.3 602 13.4
4 288 10.6 226 12.6 514 11.4
@@
Country Denmark 2,715 100.0 0 0.0 2,715 60.3
Germany 0 0.0 1,786 100.0 1,786 39.7
all 2,715 100.0 1,786 100.0 4,501 100.0
@@ A.3 Summary statistics for the consumption and price outcomes during the pre-tax and tax period
Product Outcome Pre-tax DK Tax DK $\Delta_{\text{DK}}$ Pre-tax DE Tax DE $\Delta_{\text{DE}}$ Bacon Weight 562.67 486.51 -76.16 237.36 240.64 3.28 Amount 1.89 1.80 -0.09 1.07 1.11 0.04 Expenditure 359.84 331.98 -27.86 157.18 169 11.82 Price 67.74 71.00 3.26 77.45 82.93 5.48 Butter Weight 1,607.62 1,554.70 -52.92 1,507.32 1,483.01 -24.31 Amount 6.27 6.22 -0.05 6.13 6.00 -0.13 Expenditure 913.97 1,068.94 154.97 595.71 625.36 29.65 Price 59.65 72.36 12.71 41.06 43.41 2.35 Cheese Weight 2,615.39 2,059.62 -555.77 2,678.16 2,569.77 -108.39 Amount 3.96 3.52 -0.44 9.44 9.30 -0.14 Expenditure 2,236.91 1,845.75 -391.16 1,942.41 1,948.03 5.62 Price 91.83 97.02 5.19 74.37 78.13 3.76 Cream Weight 1,465.81 1,548.64 82.83 1,099.53 1,108.92 9.39 Amount 3.47 3.53 0.06 5.10 5.14 0.04 Expenditure 481.91 569.35 87.44 242.29 267.57 25.28 Price 34.95 39.24 4.29 23.03 25.14 2.11 Liver sausage Weight 1,140.82 980.60 -160.22 95.21 88.36 -6.85 Amount 3.32 2.84 -0.48 0.66 0.59 -0.07 Expenditure 615.14 532.67 -82.47 78.02 74.96 -3.06 Price 61.48 63.19 1.71 86.28 91.05 4.77 Margarine Weight 1,842.00 1,693.05 -148.95 1,969.10 1,760.18 -208.92 Amount 3.87 3.63 -0.24 4.12 3.58 -0.54 Expenditure 511.17 570.55 59.38 368.39 350.34 -18.05 Price 29.35 34.78 5.43 19.87 21.29 1.42 Salami Weight 398.79 338.49 -60.3 474.31 463.43 -10.88 Amount 1.93 1.81 -0.12 3.05 2.97 -0.08 Expenditure 478.57 405.15 -73.42 449.1 452.91 3.81 Price 132.91 133.01 0.1 111.46 114.04 2.58 Sour cream Weight 613.76 604.32 -9.44 506.46 524.83 18.37 Amount 1.33 1.30 -0.03 2.47 2.55 0.08 Expenditure 176.17 186.28 10.11 127.57 132.27 4.70 Price 30.57 32.43 1.86 26.01 25.96 -0.05 DK indicates Danish households and DE indicates Northern-German households. $\Delta$ indicates the difference.
Product Outcome Pre-tax DK Tax DK $\Delta_{\text{DK}}$ Pre-tax DE Tax DE $\Delta_{\text{DE}}$ Eggs Weight 42.67 41.14 -1.53 41.95 39.66 -2.29 Amount 3.70 3.65 -0.05 4.41 4.20 -0.21 Expenditure 960.49 963.13 2.64 697.3 641.03 -56.27 Price 2,538.28 2,582.03 43.75 1,764.54 1,758.29 -6.25 Fruits Weight 2,056.73 730.00 -1,326.73 18,663.45 16,788.45 -1,875.00 Amount 17.36 6.28 -11.08 18.15 17.21 -0.94 Expenditure 3,634.88 1,244.02 -2,390.86 2,921.78 2,764.65 -157.13 Price 1,125.04 1,203.36 78.32 17.77 17.49 -0.28 Milk Weight 27,584.82 26,717.73 -867.09 21,434.25 18,603.68 -2,830.57 Amount 27.58 26.83 -0.75 21.10 18.39 -2.71 Expenditure 2,230.4 2,218.49 -11.91 1,200.08 1,065.29 -134.79 Price 8.67 8.91 0.24 5.84 6.00 0.16 Roast beef Weight 138.53 69.99 -68.54 31.20 46.35 15.15 Amount 0.89 0.80 -0.09 0.31 0.43 0.12 Expenditure 251.28 149.48 -101.80 98.14 126.28 28.14 Price 229.07 233.34 4.27 317.64 283.44 -34.20 Toilet paper weight 19.57 18.53 -1.04 18.61 18.44 -0.17 Amount 2.26 2.27 0.01 1.90 1.90 0.00 Expenditure 580.54 564.47 -16.07 503.12 503.69 0.57 Price 3,107.14 3,136.17 29.03 2,721.62 2,742.86 21.24 Yoghurt Weight 6,397.78 6,285.79 -111.99 4,963.91 5,062.14 98.23 Amount 7.31 7.25 -0.06 18.67 19.05 0.38 Expenditure 1,092.00 1,108.03 16.03 894.50 896.38 1.88 Price 18.53 19.14 0.61 18.93 19.08 0.15 DK indicates Danish households and DE indicates Northern-German households. $\Delta$ indicates the difference.
@@ Appendix B Fitted propensity scores for overlap verification
Appendix C Tables Anticipation Effects and Two-way Fixed Effects
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml $ATT_{2011-Q2}$20.188 -17.318 -115.138 -19.246 4.789 -57.998 60.493 28.426 (se) (27.741) (56.886) (134.277) (27.988) (24.674) (29.855) (45.022) (13.440) p-value 0.692 0.810 0.547 0.643 0.894 0.165 0.358 0.183 $ATT_{2011-Q3}$ 18.208 35.762 -78.247 -10.580 27.649 53.792 91.394 20.983 (se) (42.933) (100.254) (143.327) (87.195) (63.777) (135.263) (49.936) (48.075) p-value 0.656 0.667 0.691 0.799 0.457 0.223 0.170 0.301 Panel B: Quarterly average expenditure in Euro cents $ATT_{2011-Q2}$0.034 0.220 -0.133 -0.335 -0.005 -0.149 0.311 0.274 (se) (0.211) (0.411) (0.495) (0.348) (0.150) (0.261) (0.269) (0.207) $ATT_{2011-Q3}$ 0.227 0.112 -0.384 -0.259 0.177 0.509 0.565 0.205 (se) (0.173) (0.380) (0.479) (0.312) (0.137) (0.291) (0.279) (0.183) p-value 0.357 0.846 0.576 0.567 0.378 0.221 0.167 0.436 Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors are clustered at household level. The results are obtained controlling for the age of the head of the household, the number of children under 15 years, the gender of diary keeper, the educational level (ISCED), the income level, and the household size.
Table 11: TWFE estimates.
(1) (2) (3) (4) (5) (6) (7) (8) Bacon Butter Cheese Cream Liver sausage Margarine Salami Sour cream Panel A: Quarterly average weight consumed in gr/ml Danish fat tax (ATT) $-80.801^{***}$ $-31.697^{**}$ $-446.679^{***}$ $68.062^{**}$ $-153.833^{***}$ $57.257^{**}$ $-49.589^{***}$ $-28.427^{**}$ (se) (0.186) (0.605) (0.330) (4.545) (0.466) (0.941) (0.184) (0.473) p-value 0.001 0.012 0.000 0.042 0.002 0.010 0.002 0.011 % change -13.05% -2.04% -16.07% 5.03% -13.29% 2.98% -12.39% -5.16% Panel B: Quarterly average expenditure in Euro cents Danish fat tax (ATT) $-40.623^{***}$ $123.457^{***}$ $-396.318^{***}$ $61.382^{***}$ $-79.730^{***}$ $77.253^{***}$ $-77.510^{***}$ $5.250^{**}$ (se) (0.148) (0.176) (0.066) (0.758) (0.215) (0.030) (0.125) (0.097) p-value 0.002 0.001 0.000 0.008 0.002 0.000 0.001 0.012 % change -10.39% 13.88% -16.71% 13.33% -12.72% 14.37% -16.29% 3.30% Significance level: * $p<0.1$* $p<0.05$** $p<0.01$. Standard errors clustered at unit level.
Appendix D Evolution of tax effects over time
D.1 Bacon